If (an)infinityn=1 is a sequence of real numbers satisfying an+12 =2an-1, then,

limn->infinity an = ?


1
Expert's answer
2020-11-29T17:33:53-0500

Let "a_1=p," so

"a_2=\sqrt{2p-1},"

"a_3=\sqrt{2\sqrt{2p-1}-1},"

"a_4=\sqrt{2\sqrt{2\sqrt{2p-1}-1}-1}, ...," (can be seen, that "p\geqslant 1")

"a_k\sim (2p)^{\frac{1}{k-1}}."

So "\underset{n→\infin}{lim}a_n=\underset{n→\infin}{lim}(2p)^{\frac{1}{n-1}}=\underset{n→\infin}{lim}2^{\frac{1}{n-1}}p^{\frac{1}{n-1}}=1\cdot1=1."